# Sparse-PGD: A Unified Framework for Sparse Adversarial Perturbations Generation

### v3

@@ 6.6 Transferability of Adversarial Perturbations
To evaluate the transferability of our attack across different models and architectures, we generate adversarial perturbations based on one model and report the attack success rate (ASR) on another model. As shown in Table 6, sPGD exhibits better transferability than the most competitive baseline, Sparse-RS (RS) [17]. This further demonstrates the effectiveness of our method and its potential application in more practical scenarios.
6.7 Visualization
@@
We show some adversarial examples with different types of sparse perturbations in Figure 5, including $l_{0}$ bounded perturbations, row-wise perturbations, star-like perturbations, and adversarial watermarks of letter “A”. The attack is sPGD, and the model is vanilla ResNet-34 trained on ImageNet-100. More adversarial examples are shown in Appendix D.
7 Conclusion
@@
Appendix A Implementation Details
In experiments, we mainly focus on the cases of the sparsity of perturbations $\epsilon=10,15$ and $20$, where $\epsilon=||\sum_{i=1}^{c}\bm{\delta}^{(i)}||_{0}$ or $||{\bm{m}}||_{0}$, $\bm{\delta}^{(i)}\in{\mathbb{R}}^{h\times w}$ is the $i$-th channel of perturbation $\bm{\delta}\in{\mathbb{R}}^{h\times w\times c}$, and ${\bm{m}}\in{\mathbb{R}}^{h\times w\times 1}$ is the sparsity mask in the decomposition of $\bm{\delta}={\bm{p}}\odot{\bm{m}}$, ${\bm{p}}\in{\mathbb{R}}^{h\times w\times c}$ denotes the magnitude of perturbations.
To exploit the strength of these attacks in reasonable running time, we run all these attacks for either $10000$ iterations or the number of iterations where their performances converge. More details are elaborated below.
CornerSearch [13]: For CornerSearch, we set the hyperparameters as following: $N=100,N_{iter}=3000$, where $N$ is the sample size of the one-pixel perturbations, $N_{iter}$ is the number of queries. For bot CIFAR-10 and CIFAR-100 datasets, we evaluate the robust accuracy on $1000$ test instances due to its prohibitively high computational complexity.
Sparse-RS [17]: For Sparse-RS, we set $\alpha_{\text{init}}=0.8$, which controls the set of pixels changed in each iteration. Cross-entropy loss is adopted. Following the default setting in [17], we report the results of untargeted attacks with the maximum queries up to $10000$.
SparseFool [12]: We apply SparseFool following the official implementation and use the default value of the sparsity parameter $\lambda=3$. The maximum iterations per sample is set to $3000$. Finally, the perturbation generated by SparseFool is projected to the $l_{0}$ ball to satisfy the adversarial budget.
PGD0 [13]: For PGD0, we include both untargeted attack and targeted attacks on the top-$9$ incorrect classes with the highest confidence scores. We set the step size to $\eta=120000/255$. Contrary to the default setting, the iteration numbers of each attack increase from $20$ to $300$. Besides, $5$ restarts are adopted to boost the performance further.
SAIF [41]: Similar to $\mathrm{PGD_{0}}$, we apply both untargeted attack and targeted attacks on the top-$9$ incorrect classes with $300$ iterations per attack, however, the query budget is only $100$ iterations in the original paper [41]. We adopt the same $l_{\infty}$ norm constraint for the magnitude tensor $p$ as in sPGD.
Sparse-PGD (sPGD): Cross-entropy loss is adopted as the loss function of both untargeted and targeted versions of our method. The small constant $\gamma$ to avoid numerical error is set to $2\times 10^{-8}$. The number of iterations $T$ is $10000$ for all datasets to ensure fair comparison among attacks in Table 1. For generating $l_{0}$ bounded perturbations, the step size for magnitude ${\bm{p}}$ is set $\alpha=0.25\times\epsilon_{\infty}$; the step size for continuous mask $\widetilde{{\bm{m}}}$ is set $\beta=0.25\times\sqrt{h\times w}$, where $h$ and $w$ are the height and width of the input image ${\bm{x}}\in{\mathbb{R}}^{h\times w\times c}$, respectively; the tolerance for reinitialization $t$ is set to $3$. For generating structured sparse perturbations, we let $\alpha=0.0125\times\epsilon_{\infty},~{}\beta=0.0125\times\sqrt{h\times w}$ and $t=50$.
Sparse-AutoAttack (sAA): It is a cascade ensemble of five different attacks, i.e., a) untargeted sPGD with unprojected gradient (sPGDu), b) untargeted sPGD with sparse gradient (sPGDp), and c) untargeted Sparse-RS. The hyper-parameters of sPGD are the same as those listed in the last paragraph.
Adversarial Training: sPGD is adopted as the attack during the training phase, the number of iterations is $20$, and the backward function is randomly selected from the two different backward functions for each batch. For sTRADES, we only compute the TRADES loss when training, and generating adversarial examples is based on cross-entropy loss. We use PreactResNet18 [68] with softplus activation [71] for experiments on CIFAR-10 and CIFAR-100 [64], and ResNet34 [67] for experiments on ImageNet100 [65] and GTSRB [66]. We train the model for $100$ epochs on CIFAR-10 and CIFAR-100, for $40$ epochs on ImageNet100, and for $20$ epochs on GTSRB. The training batch size is $128$ on CIFAR-10 and CIFAR-100, and $32$ on ImageNet100 and GTSRB. The optimizer is SGD with a momentum factor of $0.9$ and weight decay factor of $5\times 10^{-4}$. The learning rate is initialized to $0.05$ and is divided by a factor of $10$ at the $\frac{1}{4}$ and the $\frac{3}{4}$ of the total epochs. The tolerance for reinitialization is set to $10$.
Appendix B More Results of Section 6.1
In this subsection, we present the robust accuracy on CIFAR-10 with the sparsity level $\epsilon=10$ and $\epsilon=15$, as well as those on CIFAR-100 with $\epsilon=10$ in Table 7, 8 and 9, respectively. The observations with different sparsity levels and on different datasets are consistent with those in Table 1 and 2, which indicates the effectiveness of our method.
Table 7: Robust accuracy of various models on different sparse attacks, where the sparsity level $\epsilon=10$. The models are trained on CIFAR-10 [64]. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in [17]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
Vanilla RN-18 93.9 3.2 0.5 40.6 11.5 31.8 0.5 7.7 0.5
$l_{\infty}$-adv. trained, $\epsilon=8/255$
GD PRN-18 87.4 36.8 24.5 69.9 50.3 63.0 31.0 37.3 23.2
PORT RN-18 84.6 36.7 27.5 70.7 46.1 62.6 31.0 36.2 25.0
DKL WRN-28 92.2 40.9 25.0 71.9 54.2 64.6 32.6 38.8 23.7
DM WRN-28 92.4 38.7 23.7 68.7 52.7 62.5 31.2 37.4 22.6
$l_{2}$-adv. trained, $\epsilon=0.5$
HAT PRN-18 90.6 47.3 40.4 74.6 53.5 71.4 37.3 36.7 34.5
PORT RN-18 89.8 46.8 37.7 74.2 50.4 70.9 33.7 33.0 30.6
DM WRN-28 95.2 57.8 47.9 78.3 65.5 80.9 47.1 48.2 43.1
FDA WRN-28 91.8 55.0 49.4 79.6 58.6 77.5 46.7 49.0 43.8
$l_{1}$-adv. trained, $\epsilon=12$
$l_{1}$-APGD PRN-18 80.7 51.4 51.1 74.3 60.7 68.1 47.2 47.4 45.9
Fast-EG-$l_{1}$ PRN-18 76.2 49.7 48.0 69.7 56.7 63.2 44.7 45.0 43.2
$l_{0}$-adv. trained, $\epsilon=10$
PGD0-A PRN-18 85.8 20.7 16.1 77.1 66.1 68.7 33.5 36.2 15.1
PGD0-T PRN-18 90.6 22.1 14.0 85.2 72.1 76.6 37.9 44.6 13.9
sAT PRN-18 86.4 61.0 57.4 84.1 82.0 81.1 78.2 77.6 57.4
sTRADES PRN-18 89.8 74.7 71.6 88.8 87.7 86.9 85.9 84.5 71.6
Table 8: Robust accuracy of various models on different sparse attacks, where the sparsity level $\epsilon=15$. The models are trained on CIFAR-10 [64]. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in [17]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
Vanilla RN-18 93.9 1.6 0.0 25.3 2.1 12.0 0.0 0.0 0.0
$l_{\infty}$-adv. trained, $\epsilon=8/255$
GD PRN-18 87.4 30.5 12.2 61.1 36.0 51.3 17.1 24.3 11.3
PORT RN-18 84.6 30.8 15.2 62.1 31.4 52.1 17.4 23.0 13.0
DKL WRN-28 92.2 35.3 13.2 62.5 41.2 52.3 18.4 24.9 12.1
DM WRN-28 92.4 34.8 12.6 57.9 38.5 49.4 17.9 24.0 11.6
$l_{2}$-adv. trained, $\epsilon=0.5$
HAT PRN-18 90.6 38.9 23.5 65.3 35.4 60.2 19.0 18.6 16.6
PORT RN-18 89.8 36.8 20.6 64.3 30.6 59.7 16.0 15.5 13.8
DM WRN-28 95.2 48.5 27.7 68.2 47.5 70.9 25.0 26.7 22.1
FDA WRN-28 91.8 47.8 31.1 71.8 40.1 68.2 28.0 31.4 25.5
$l_{1}$-adv. trained, $\epsilon=12$
$l_{1}$-APGD PRN-18 80.7 41.3 36.5 70.3 50.5 62.3 30.4 31.3 29.0
Fast-EG-$l_{1}$ PRN-18 76.2 40.7 34.8 64.9 46.7 56.9 29.6 30.1 28.0
$l_{0}$-adv. trained, $\epsilon=15$
PGD0-A PRN-18 83.7 17.5 6.1 73.7 62.9 60.5 19.4 27.5 5.6
PGD0-T PRN-18 90.5 19.5 7.2 85.5 63.6 69.8 31.4 41.2 7.1
sAT PRN-18 80.9 46.0 37.6 77.1 74.1 72.3 71.2 70.3 37.6
sTRADES PRN-18 90.3 71.7 63.7 89.5 88.1 86.5 85.9 83.8 63.7
Table 9: Robust accuracy of various models on different sparse attacks, where the sparsity level $\epsilon=10$. The models are trained on CIFAR-100 [64]. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in [17]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
Vanilla RN-18 74.3 1.6 0.3 20.1 1.9 9.0 0.1 0.9 0.1
$l_{\infty}$-adv. trained, $\epsilon=8/255$
HAT PRN-18 61.5 12.6 9.3 39.1 19.1 26.8 11.6 14.2 8.5
FDA PRN-18 56.9 16.3 12.3 42.2 23.0 30.7 14.9 17.8 11.6
DKL WRN-28 73.8 12.4 6.3 44.9 20.9 26.5 10.5 14.0 6.1
DM WRN-28 72.6 14.0 8.2 46.2 23.4 29.8 12.7 15.8 8.0
$l_{1}$-adv. trained, $\epsilon=6$
$l_{1}$-APGD PRN-18 63.2 22.7 22.1 47.7 33.0 43.5 19.7 20.3 18.5
Fast-EG-$l_{1}$ PRN-18 59.4 21.5 21.0 44.8 30.6 39.5 18.9 18.6 17.3
$l_{0}$-adv. trained, $\epsilon=10$
PGD0-A PRN-18 66.1 9.3 7.1 57.9 29.9 39.5 13.9 20.4 6.5
PGD0-T PRN-18 70.7 14.8 10.5 63.5 46.3 51.7 24.5 28.6 10.2
sAT PRN-18 67.0 44.3 41.6 65.9 61.6 60.9 56.8 58.0 41.6
sTRADES PRN-18 70.9 52.8 50.3 69.2 67.2 65.2 64.0 63.7 50.2
Appendix C Ablation Studies
We conduct ablation studies in this section. We focus on CIFAR10 and the sparsity level $\epsilon=20$. Unless specified, we use the same configurations as in Table 1.
Table 10: Ablation study of each component in sPGDproj in terms of robust accuracy. The model is Fast-EG-$l_{1}$ trained on CIFAR-10.
Ablations Robust Acc.
Baseline (PGD0 w/o restart) 49.4
    + Decomposition: $\bm{\delta}={\bm{p}}\odot{\bm{m}}$ 58.0 (+8.6)
    + Continuous mask $\widetilde{{\bm{m}}}$ 33.9 (-15.5)
    + Random reinitialization 18.1 (-31.3)
sPGD: We first validate the effectiveness of each component in sPGD. The result is reported in Table 10. We observe that naively decomposing the perturbation $\bm{\delta}$ by $\bm{\delta}={\bm{p}}\odot{\bm{m}}$ and updating them separately can deteriorate the performance. By contrast, the performance significantly improves when we update the mask ${\bm{m}}$ by its continuous alternative $\widetilde{{\bm{m}}}$ and $l_{0}$ ball projection. This indicates that introducing $\widetilde{{\bm{m}}}$ greatly mitigates the challenges in optimizing discrete variables. Moreover, the results in Table 10 indicate the performance can be further improved by the random reinitialization mechanism, which encourages exploration and avoids trapping in a local optimum.
In addition, we compare the performance when we use different step sizes for the magnitude tensor ${\bm{p}}$ and the sparsity mask ${\bm{m}}$. As shown in Table D and D, the robust accuracy does not vary significantly with different step sizes. It indicates the satisfying robustness of our method to different hyperparameter choices. In practice, We set $\alpha$ and $\beta$ to $0.25$ and $0.25\times\sqrt{hw}$, respectively. Note that $h$ and $w$ denote the height and width of the image, respectively, which are both $32$ in CIFAR-10.
Ultimately, we compare the performance of sPGD with different tolerance iterations for reinitialization. As shown in Table D, the performance of our method remains virtually unchanged, which showcases that our approach is robust to different choices of tolerance for reinitialization.
Adversarial training: We conduct preliminary exploration on adversarial training against sparse perturbations, since sPGD can be incorporated into any adversarial training variant. Table 1278 and 9 study sAT and sTRADES, while sTRADES outperforms sAT in all cases. In addition, the training of sAT is relatively unstable in practice, so we focus on sTRADES for ablation studies in this section. We leave the design of sPGD-adapted adversarial training variants to further improve model robustness as a future work.
Table D demonstrates the performance when we use different backward functions. The policies include always using the sparse gradient (Proj.), always using the unprojected gradient (Unproj.), alternatively using both backward functions every $5$ epochs (Alter.) and randomly selecting backward functions (Rand.). The results indicate that randomly selecting backward functions has the best performance. In addition, Table D demonstrates the robust accuracy of models trained by sTRADES with different multi-$\epsilon$ strategies. The results indicate that multi-$\epsilon$ strategy helps boost the performance. The best robust accuracy is obtained when the adversarial budget for training is $6$ times larger than that for test. Furthermore, we also study the impact of different tolerances during adversarial training in Table D. The results show that higher tolerance during adversarial training benefits the robustness of the model, and the performance reaches its best when the tolerance is set to $10$.
Structured Sparse Perturbations: Since the optimal hyperparameters for structured sparse perturbations differ from those for unstructured $l_{0}$ bounded perturbations, we conduct further ablation studies in the structured sparse setting.
The results in Table D, D and D suggest that the step size $\alpha$, $\beta$ and the tolerance $t$ for generating structured sparse perturbations should be smaller than those for unstructured cases. This can be attributed to the requirement for more exploitable attacks when generating structured sparse perturbations, similar to conventional $l_{\infty}$ and $l_{2}$ attacks [39].
The multi-$\epsilon$ strategy has been proven effective in $l_{0}$ adversarial training. However, in the context of structured sparse attack, particularly in the patch setting, we can increase either the sparsity level $\epsilon$ or the patch size to increase the corresponding $l_{0}$ adversarial budget. Therefore, we compare two different strategies here, i.e., multi-$\epsilon$ strategy and multi-size strategy. Specifically, the multi-$\epsilon$ strategy increases the sparsity level $\epsilon$ while maintaining the patch size unchanged during training, but the multi-size strategy only increases the patch size to make the $l_{0}$ sparsity level of generated perturbations similar to that in the multi-$\epsilon$ strategy. The results reported in Table D indicate that the multi-$\epsilon$ strategy provides better performance.
Appendix D More Visualization
More adversarial examples are shown in Figure 6. Additionally, we showcase some $l_{0}$ bounded adversarial examples generated on our sAT model in Figure 7. Compared to Figure 5 (a)-(b) and Figure 6 (a)-(b), the perturbations generated on the adversarially model are mostly located in the foreground of images. It is consistent with the intuition that the foreground of an image contains most of the semantic information [41].
Table 11: Robust accuracy at different step sizes $\alpha$ for magnitude $p$. The evaluated attack is sPGDproj. The model is Fast-EG-$l_{1}$ trained on CIFAR-10.
$\alpha$ $\frac{1}{16}$ $\frac{1}{8}$ $\frac{1}{4}$ $\frac{1}{2}$ $\frac{3}{4}$ $1$
Acc. 19.6 18.8 18.1 18.4 18.7 19.0
Table 12: Robust accuracy at different step sizes $\beta$ for mask $m$. The evaluated attack is sPGDproj. The model is Fast-EG-$l_{1}$ trained on CIFAR-10. Note that $h$ and $w$ are both $32$ in CIFAR-10.
$\beta/\sqrt{hw}$ $\frac{1}{16}$ $\frac{1}{8}$ $\frac{1}{4}$ $\frac{1}{2}$ $\frac{3}{4}$ $1$
Acc. 20.0 19.3 18.1 18.4 19.4 21.0
Table 13: Robust accuracy at different tolerance for reinitialization $t$ during attacking. The evaluated attack is sPGDproj. The model is Fast-EG-$l_{1}$ trained on CIFAR-10.
$t$ 1 3 5 7 10
Acc. 18.1 18.1 18.5 18.5 18.5
Table 14: Ablation study on different policies during adversarial training. The model is PRN18 trained by sTRADES with $6\times k$. The robust accuracy is obtained through sAA.
Policy Proj. Unproj. Alter. Rand.
Acc. 41.5 39.4 51.5 61.7
Table 15: Ablation study on multi-$\epsilon$ strategy during adversarial training. The model is PRN18 trained by sTRADES with random policy. The robust accuracy is obtained through sAA.
$\epsilon$ 1×1$\times$1 × 2×2$\times$2 × 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 × 10×10$\times$10 ×
Acc. 34.0 39.7 54.5 61.7 60.2 55.5
Table 16: Ablation study on tolerance for reinitialization $t$ during adversarial training. The model is PRN18 trained by sTRADES with $6\times k$ and tolerance $t=3$. The robust accuracy is obtained through sAA.
$t$ 3 10 20
Acc. 51.7 61.7 60.0
Table 17: Robust accuracy at different step sizes $\alpha$ for magnitude $p$. The evaluated attack is sPGDproj generating $5\times 5$ patch ($\epsilon=1$). The model is sTRADES-$l_{0}$trained on CIFAR-10.
$\alpha$ $0.01$ $0.0125$ $0.05$ $0.1$ $0.15$ $0.25$
Acc. 28.1 26.7 33.8 41.5 47.9 58.1
Table 18: Robust accuracy at different step sizes $\beta$ for mask $m$. The evaluated attack is sPGDproj generating $5\times 5$ patch ($\epsilon=1$). The model is sTRADES-$l_{0}$ trained on CIFAR-10. Note that $h$ and $w$ are both $32$ in CIFAR-10.
$\beta/\sqrt{hw}$ $0.01$ $0.0125$ $0.05$ $0.1$ $0.15$ $0.25$
Acc. 27.4 26.7 34.1 37.5 38.3 35.1
Table 19: Robust accuracy at different tolerance for reinitialization $t$ during attacking. The evaluated attack is sPGDproj generating $5\times 5$ patch ($\epsilon=1$). The model is sTRADES-$l_{0}$ trained on CIFAR-10.
$t$ 3 10 50 100 150
Acc. 36.1 28.7 26.7 28.1 33.7
Table 20: Ablation study on multi-$\epsilon$ / multi-size strategy during adversarial training against $3\times 3$ patch ($\epsilon=2$). The model is PRN18 trained by sTRADES with random policy.
Attack RS LOAP sPGDproj sPGDunproj sAA
Multi-$\epsilon$ 77.5 81.4 72.8 81.7 72.3
Multi-size 75.5 79.2 68.4 79.8 67.9
(a) $l_{0}$, $\epsilon=100$
(b) $l_{0}$, $\epsilon=100$
(c) column, $\epsilon=1$
(d) column, $\epsilon=1$
(e) patch, $14\times 14$, $\epsilon=1$
(f) patch, $14\times 14$, $\epsilon=1$
(g) patch, $10\times 10$, $\epsilon=2$
(h) patch, $10\times 10$, $\epsilon=2$
(i) heart, $20\times 20$, $\epsilon=1$
(j) heart, $20\times 20$, $\epsilon=1$
(k) circle, $60\times 60$, $\epsilon_{\infty}=64/255$
(l) circle, $60\times 60$, $\epsilon_{\infty}=64/255$
[…]
(a) $l_{0}$, $\epsilon=200$
(b) $l_{0}$, $\epsilon=200$
(c) $l_{0}$, $\epsilon=200$
(d) $l_{0}$, $\epsilon=200$
[…]

### v4

@@ VI-F Transferability of Adversarial Perturbations
To evaluate the transferability of our attack across different models and architectures, we generate adversarial perturbations based on one model and report the attack success rate (ASR) on another model. As shown in Table VI, sPGD exhibits better transferability than the most competitive baseline, Sparse-RS (RS) [17]. This further demonstrates the effectiveness of our method and its potential application in more practical scenarios.
VI-G Attack Object Detection and Segmentation Models
The versatility of our method allows it to be applied to a variety of tasks beyond classification, e.g., object detection and segmentation. For implementation, one could simply replace the loss function $\mathcal{L}$ in Algorithms 1 and 2 with a specific detection or segmentation loss function.
In Table VII, we report the mAP50-95 of YOLOv8-M [70] in object detection and segmentation tasks when the input image is manipulated by sparse perturbations, including both structured and unstructured ones. In addition, we include the results of RS and SAIF as a comparison. Since SAIF is not able to generate structured sparse perturbations, we do not include its results in the experiments of generating patch perturbations. The results indicate that sPGD, as an attack framework, is still effective in these two tasks. Considering higher resolution images in the COCO dataset [71] ($640\times 640$), we use approximately the same ratio between the number of perturbed pixels and the total number of pixels as in Table II when setting the perturbation budgets $\epsilon$, i.e., $0.4\%$ / $0.8\%$ of the pixels. Compared with structured perturbations, unstructured perturbations are spatially more dispersed, making them more effective for multi-object images in these tasks.
TABLE VII: Evaluation of YOLOv8-M performance on clean and adversarial COCO dataset [71] in object detection and segmentation tasks. The metric is mAP50-95. We test the performance on the adversarial examples generated by different attacks, where both unstructured and structured cases are considered.
Mode $\epsilon$ Clean RS SAIF sPGD
$l_{0}$ 1600 55.1 17.1 15.8 15.5
3200 55.1 8.1 7.4 6.6
$4\times 4$ patch 100 55.1 47.1 - 40.1
200 55.1 40.8 - 34.5
(a) Object Detection
Mode $\epsilon$ Clean RS SAIF sPGD
$l_{0}$ 1600 53.6 17.6 16.4 15.6
3200 53.6 8.8 8.2 7.4
$4\times 4$ patch 100 53.6 46.5 - 39.1
200 53.6 39.7 - 33.1
(b) Segmentation
VI-H Visualization
@@
For a better visualization, we present some adversarial examples with different types of sparse perturbations in Figure 5, including $l_{0}$ bounded perturbations, row-wise perturbations, star-like perturbations, and adversarial watermarks of letter “A”. The attack is sPGD, and the model is vanilla ResNet-34 trained on ImageNet-100. More adversarial examples are shown in Appendix D.
VII Conclusion
@@
Acknowledgments
This work is supported by the NSFC project (No. 62306250). It is supported by the internal funds of City University of Hong Kong (No. 9610614 and No. 9229130). We also thank Yixiao Huang for his contribution to this work.